6) A 39.0kg block at rest on a horizontal frictionless table is connected to the wall via a spring with a spring constant k=16.0N/m. A 3.70
Answer #1

initial momentum = Pi = m*v

final momentumPf = (M+m)*V

Pf = Pi

V = mv/(M+m)

KE = PE

0.5*(M+m)*V^2 = 0.5*K*A^2

(0.5*0.037^2*560^2)/(39+0.037) = 0.5*16*A^2

A = 0.8291 m


7) F = P * A = (Po+rho*g*h)*A = ((1.013*10^5)+1024)*9.8*113*3.14*(0.171/2)^2 = 2.601*10^6 N